1. 50.00 mL of 0.1000 M propanoic acid (CH3CH2COOH – Ka = 1.34 X 10-5) is titrated with 0.2000 M KOH. Calculate the pH at the following points in the titration: 1) Initial pH – no KOH has been added. 2) 5.00 mL of KOH has been added. 3) 12.50 mL of KOH has been added. 4) At the equivalence point. (Calculate the volume of KOH to reach the equivalence point & identify a good indicator.) 5) Provide a sketch of the titration showing the volume & pH at significant points.
Answer #1

1.) [CH3CH2COOH]=0.1000 M and Ka = 1.34 X 10-5.

ka=[CH3CH2COO-][H+]/[CH3CH2COOH]=[H+]2/0.10=1.34 X 10-5.

So, [H+]=(1.34 X 10-6)1/2=1.158*10-3. pH=-log10[H+]=2.936.

2.) 50.00 mL of 0.1000 M propanoic acid=5/1000 mole=5 mmole.

5.00 mL of 0.2000 M NaOH=1/1000 mole=1 mmole.

1 mmole NaOH reacts with 1 mmole propanoic acid to form 1 mmole salt (CH3CH2COONa) of the acid. Total volume after addition of 5.00 mL KOH is 50+5=55 ml. Moles of acid remains unreacted is (5-1) mmole=4 mmole.

[CH3CH2COOH]=4 mmole*1000 ml/55 ml=0.0727 M.

[CH3CH2COONa]=1 mmole*1000 ml/55 ml=0.0182 M.

Let's take, [H+]=x M.

         CH3CH2COOH <-> H+ +CH3CH2COO-

I            0.0727                        0.0182

C             -x                     x          +x

E          0.0727-x             x        0.0182+x

ka=[CH3CH2COO-] [H+]/[CH3CH2COOH]=(0.0182+x)x/( 0.0727-x)=1.34*10-5. So, x=0.5353*10-4.

pH=-log10[H+]=4.27

3.) Similar to above calculation, one can have

[CH3CH2COOH]=2.5 mmole*1000 ml/62.5 ml=0.04 M.

[CH3CH2COONa]=2.5 mmole*1000 ml/62.5 ml=0.04 M.

x~1.34*10-5. pH=-log10[x]=4.87

4.) At equivalance point, 5 mmole acid reacts with 5mmole base. So the required volume of NaOH is 25 ml.

The total volume of the reaction mixture is 50+25=75 ml.

So, the concentration of salt is 5 mmole/75 ml=0.0667 M.

CH3CH2COO-+H2O <-> CH3CH2COOH+OH-

kb=kw/ka=10-14/1.34*10-5=[OH-][CH3CH2COOH]/[CH3CH2COO-]=[OH-]2/0.0667 M.

So, [OH-]=0.223*10-4. pH=14-pOH=14+log10[OH-]=9.35